INFRARED SPECTROSCOPY OF LARGE-SIZED PHENOL-WATER CLUSTERS PhOH-(H_{2}O)_{n} ( 10 ≤ n ≤ 50)

\underline{TORU HAMASHIMA}, KENTA MIZUSE and ASUKA FUJII, Department of Chemistry, Graduate School of Science, Tohoku University, Sendai 980-8578,Japan.

We report infrared spectra of moderately size-selected phenol-(H_{2}O)_{n−1} ( 10 ≤ n ≤ 50), which have essentially the same network structures as (H_{2}O)_{n}. The spectra in the OH stretching region are observed. Detailed analyses of these spectra aided by density functional theory calculations reveal the development process of the hydrogen bond network.